OpenStudy (anonymous):

heres a new one....write the equation of a line in slope-intercept form with the given slope and point P: Slope = 2; P(7,8)

jigglypuff314 (jigglypuff314):

slope-intercept form: y = mx + b m = slope = 2 so you get y = 2x + b then you can temporarily plug in the given point to find b so when given the point (7, 8) you plug in and get 8 = 2(7) + b solve for b and once you get b plug in what you got for b into y = 2x + b

OpenStudy (anonymous):

so y=2x+(-6) ?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

What if you had..Write the equation of the line (in slope-intercept form) passing through P and Q: P(3,7) Q(2,12)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ P&(3\quad ,&7)\quad Q&(2\quad ,&12) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

This equation is that of a function: y = 2x+3...... true or false?